OpenStudy (anonymous):

5y + 2 = 1/2 (10y + 4 )

OpenStudy (anonymous):

whats the 1st step you will take?

OpenStudy (anonymous):

I think it is distribute

OpenStudy (anonymous):

nice, do it, what do you get?

OpenStudy (anonymous):

wait what do i distribute..?

OpenStudy (anonymous):

a*(b+c) = a*b+a*c

OpenStudy (anonymous):

ummmm im going to try i am really bad at math

OpenStudy (anonymous):

try

OpenStudy (anonymous):

ok i did 0.5 time ten = 5 what do i do now..?

OpenStudy (anonymous):

write the whole thing so i can see what did you do

OpenStudy (anonymous):

ok 5y + 2 = 0.5 and i put 5

OpenStudy (anonymous):

its .5*(10y + 4) so what is it equal to?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

wait i already got 5 * ten

OpenStudy (anonymous):

write the whole thing

OpenStudy (anonymous):

ok what i got so far is 5y + 2 = 0.5 (10y + 4) 5y + 2 = 5y + 2 -2 -2 5y/5y 5y/5y y=1 am i right..?

OpenStudy (anonymous):

you got it did you do it alone?

OpenStudy (anonymous):

nope my sister she is yelling at me ughh

OpenStudy (anonymous):

hehe ok good